Suppose in the fall semester the student union allows anyone to sell bottled water. You observe that the P=$1 per bottle and 1000 bottles are sold per week. When you return to campus in the Spring, you notice that the student union now allows only Pepsi to sell bottled water. You observe that the price has risen to $2 per bottle and the quantity sold per week is 600 bottles. The economics department has estimated that if the price rises to $5 per bottle zero units will be sold per week. What is the loss in consumer surplus from the Fall to the Spring semester?

A. 900

B. 2,000

C. 4000

D. 1,100
Answer #1

The intercept of the budget line is 5 because at this price the quantity demanded is 0.

Initial consumer surplus=ABC when price=$1

Area of ∆ABC=(AB*BC)/2

AB=5-1=4

BC=1000

ABC=(4*1000)/2

Area∆ABC=2000

New consumer surplus at price=$2

CS= Area AB'C'

A'B'=5-2=3

B'C'=600

CS=(3*600)/2=900

Loss in CS=Original CS-New CS

Loss=2000-900=1100

Answer-D

Please rate the answer.